To find g(x+2) what do you substitute for x in the function g(x)=8x-2
Answer by Edwin McCravy(20055)

To get g(x+2) we start out with this:

g(x) = 8x-2

We put parentheses around the x on the right side:

g(x) = 8(x)-2

We take out both x's

g( ) = 8( )-2

We open the empty parentheses a little wider

g(   ) = 8(   )-2

We put x+2 in the parentheses:

g(x+2) = 8(x+2)-2

We leave the left side as it is, but 
we simplify the right side:

g(x+2) = 8x+16-2

g(x+2) = 8x+14

Edwin